Cu(s) + AgNO3(aq) ---> CuNO3(aq) + Ag(s)

How many grams of silver would be produced if you began with 25 grams of silver nitrate?
1
Expert's answer
2020-12-24T03:54:05-0500

Moles of silver nitrate taken = 25/170 = 0.147 moles


that much moles of silver will be produced = 0.147 moles of silver


which is equal to 0.147 × 107.8 g = 15.86 g